# Fact or Guesswork? Evaluating Large Language Model’s Medical Knowledge with Structured One-Hop Judgment

### v1

@@ 4.6 RQ4: What strategies can enhance the response accuracy of LLMs?
  • Documents. We utilize the knowledge base $\mathcal{D}$ constructed and described in Section 3.2 as our retrieval document corpus.
  • Implementation. We prepend the retrieved grounding documents to the input of the LLMs Ram et al. (2023); Shi et al. (2024a). Detailed prompts can be found in Appendix A.3.
The experiment outputs are listed in Table 3. The performance differentials between RAG and zero-shot approaches are indicated in parentheses, computed as RAG results minus zero-shot results. Performances improved through RAG are displayed in green with upward arrows ($\uparrow$), while performance decrements are shown in red with downward arrows ($\downarrow$).
@@ 5 Conclusion
We introduce the Medical Knowledge Judgment (MKJ) dataset which is designed to assess the ability of LLMs to internalize medical and healthcare factual knowledge by extracting and formulating medical knowledge triplets into one-hop direct judgment questions. Our experiments on the MKJ dataset and subsequent analysis indicate that LLMs still face challenges in retaining factual medical knowledge, with notable variations in performance across different semantic categories, especially when dealing with rare medical conditions. In the meanwhile, most LLMs exhibit poor calibration status with mixed overconfidence and underconfidence patterns. To address these issues, we investigate retrieval-augmented generation, and demonstrate its effectiveness in enhancing factual accuracy and reducing uncertainty in medical decision-making.
@@ A.3 Prompts
A.4 Calibration curves and uncertainty quantification

### v2

@@ 4.6 RQ4: What strategies can enhance the response accuracy of LLMs?
  • Documents. We utilize the constructed knowledge base $\mathcal{D}$ as described in Section 3.2 as our retrieval document corpus.
  • Implementation. We retrieved documents for each query according to their cosine similarities, and prepend the top-5 to the input of the LLMs following Ram et al. (2023) and Shi et al. (2024a). Detailed prompts can be found in Appendix A.3.
The experiment outputs are listed in Table 3. The performance differentials between RAG and zero-shot approaches are indicated in parentheses, computed as RAG results minus zero-shot results.
@@ 5 Conclusion
We introduce the Medical Knowledge Judgment (MKJ) dataset which is designed to assess the factuality of LLMs in medical and healthcare by extracting and formulating knowledge triplets into one-hop direct judgment questions. Our experiments on the MKJ dataset and subsequent analysis indicate that LLMs still face challenges in retaining factual knowledge, with notable variations in performances across different medical semantic types, especially for rare medical conditions. In the meanwhile, most LLMs exhibit poor calibration status with mixed overconfidence and underconfidence patterns. To address these issues, we investigate retrieval-augmented generation, and demonstrate its effectiveness in enhancing accuracy and reducing uncertainty in medical knowledge retention.
@@ A.3 Prompts
A.4 Performances with zero-shot prompting
We display the full results of LLMs with zero-shot prompting in Table 5 below.
[…]
GPT-3.5-Turbo 0.82 0.70 0.86 -0.16
GPT-4o-mini 0.84 0.74 0.87 -0.13
GPT-4o 0.85 0.64 0.92 -0.28
Claude-3-haiku 0.72 0.88 0.66 0.22
Claude-3-sonnet 0.83 0.68 0.88 -0.20
Llama-3-8B 0.71 0.51 0.78 -0.27
Llama-3.1-8B 0.68 0.51 0.74 -0.23
Llama-3.2-1B 0.65 0.12 0.82 -0.70
Llama-3.2-3B 0.67 0.42 0.75 -0.33
Ministral-8B 0.71 0.48 0.79 -0.31
Qwen2.5-0.5B 0.56 0.50 0.58 -0.08
Qwen2.5-1.5B 0.53 0.73 0.46 0.27
Qwen2.5-3B 0.68 0.16 0.86 -0.70
Phi-3-mini-4k 0.70 0.64 0.72 -0.08
Phi-3-mini-128k 0.72 0.58 0.76 -0.18
Phi-3.5-mini 0.69 0.67 0.71 -0.04
Meditron-7B 0.82 0.74 0.84 -0.10
MeLlama-13B 0.83 0.79 0.85 -0.06
Table 5: Model performances with zero-shot prompting.
A.5 Calibration Curves and Uncertainty Quantification